A simple economy produces two goods, Bread and Technical Manuals. Price and quantity data are as follows: Production and Prices in Year 1 (Base year) Quantity 110 500 Price Per Unit Product Bread Technical Manuals $2.00 $45.00 Production and Prices in Year 2 Product Bread Technical Manuals Quantity 137.50 750.00 Price Per Unit $3.00 $90.00 In Year 2, nominal GDP is equal to: S and real GDP is S (enter both responses rounded to the nearest penny).
Answer #1

To obtain the real GDP, calculate the Nominal GDP of year 2. This is done using the formula

Nominal GDP of year 2 = Sum of (current year price * Current year quantity) = (137.5*3)+(750*90) = 412.5 + 67500 = $67,912.50

Now, Real GDP would be  = Sum of (base year price * Current year quantity) = (137.5*2)+(750*45) = 275 + 33,750 =

$34,025